# Preference Poisoning Attacks on Reward Model Learning

### v1

@@ Abstract.
Learning utility, or reward, models from pairwise comparisons is a fundamental component in a number of application domains, including autonomous control, conversational agents, and recommendation systems, as part of a broad goal of aligning automated decisions with user preferences. These approaches inherently entail collecting preference information from people, with feedback often provided anonymously. Since preferences are subjective, there is no gold standard to compare against; yet, reliance of high-impact systems on preference learning creates a strong motivation for malicious actors to skew data collected in this fashion to their ends. We investigate the nature and extent of this vulnerability systematically by considering a threat model in which an attacker can flip a small subset of preference comparisons with the goal of either promoting or demoting a target outcome. First, we propose two classes of algorithmic approaches for these attacks: a principled gradient-based framework, and several variants of rank-by-distance methods. Next, we demonstrate the efficacy of best attacks in both these classes in successfully achieving malicious goals on datasets from three diverse domains: autonomous control, recommendation system, and textual prompt-response preference learning. We find that the best attacks are often highly successful, achieving in the most extreme case 100% success rate with only 0.3% of the data poisoned. However, which attack is best can vary significantly across domains, demonstrating the value of our comprehensive vulnerability analysis that involves several classes of attack algorithms. In addition, we observe that the simpler and more scalable rank-by-distance approaches are often competitive with the best, and on occasion significantly outperform gradient-based methods. Finally, we show that several state-of-the-art defenses against other classes of poisoning attacks exhibit, at best, limited efficacy in our setting.
Poisoning attacks, reward model learning
copyright: acmlicensedjournalyear: 2018doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation emai; June 03–05, 2018; Woodstock, NYisbn: 978-1-4503-XXXX-X/18/06
1. Introduction
@@
Remarkable advances in AI ranging from autonomous systems, such as self-driving cars (Badue et al., 2021), to conversational agents (Wu et al., 2023), have led to significant concerns about the possibility that the resulting technologies would come to systematically deviate from conventional norms.
Figure 1. Illustration of Poisoning Attack (Promotion).
@@
Such deviations can range from benign, for example, an occasionally unhelpful response to a prompt, to harmful, such as generating autonomous trajectories that make users (e.g., riders in self-driving cars) fear for their lives or the lives of others, to generating hate speech or inciting violence in response to conversational prompts. The fundamental concern is to ensure that the behavior of AI systems aligns with the values of the human communities impacted by these. This value alignment problem (Brown et al., 2021; Russell, 2019) has been a subject of foundational research in a broad array of domains, which in addition to autonomous control (Christiano et al., 2017) and conversational agents (Ouyang et al., 2022), include recommendation systems (Kalloori et al., 2018; Qomariyah, 2018), and, classically, elections and policymaking (Noothigattu et al., 2018). A common foundational element of value alignment, with roots in social choice and utility theory, is to learn a utility function, often referred to as reward model, from a dataset comprising human responses about their relative preferences between pairs of options (outcomes, candidates). The learned reward model can then be used in the relevant downstream tasks, such as to make personalized recommendations, or as part of the reinforcement learning from human feedback (RLHF) which has gained popularity in the design of helpful and harmless language models (Ouyang et al., 2022; Sun et al., 2023; Stiennon et al., 2020).
@@
The nature of preference elicitation necessary in reward model learning, however, opens the door to malicious tampering, as one commonly obtains such information from otherwise anonymous subjects. Moreover, the subjective nature of preferences, which on many matters represent a potential diversity of opinions, makes it very challenging, if not impossible to evaluate the veracity of responses with respect to any gold standard. Indeed, individuals themselves may well have incentives to misreport their preferences, a significant concern due to the well-known Gibbard-Satterthwaite result (Barberá, 1983). For example, some users may have their own prejudices, and would wish to see more harmful content being generated in conversation. Additionally, malicious actors may have an incentive to create fake user accounts, or compromise legitimate users, and thereby populate pairwise comparison datasets collected with preferences responses that achieve their own ends. Consequently, a crucial and under-explored issue is one of vulnerability of reward model learning from pairwise comparison data to malicious preference label poisoning. Our goal is to systematically investigate this vulnerability within the most common Bradley-Terry (BT) reward model learning framework based on maximum likelihood estimation (MLE) (Bradley and Terry, 1952; Xia, 2019) (for example, this is the typical approach in RLHF (Ouyang et al., 2022; Sun et al., 2023; Stiennon et al., 2020)).
@@
Our work is closely related to the broader literature pertaining to poisoning attacks on machine learning generally, and label-flipping attacks in particular. The focus in the existing literature has been predominantly on either classification or regression problems. However, for reward model learning, which has features of both regression (insofar as the reward model maps feature vectors to real utility values) and classification (insofar as we learn from pairwise preferences, in which labels are binary indicators of which option is preferred), little attention has been given. Moreover, the problem setting motivates attack goals that focus specifically on relative preferences over feature vectors, which are in themselves distinct from common threat models in prior literature. We show that as a consequence of these structural differences in the learning problem and threat model, common defenses against label flipping attacks on conventional supervised learning methods have only limited efficacy in our setting.
@@
There are several technical challenges in devising preference poisoning attacks on reward model learning. The first is that the nature of pairwise is discrete: either one prefers the first option, the second, or is indifferent. Consequently, gradient-based methods proposed for data poisoning, which assume that data is real-valued, cannot be used directly. On the other hand, the typical heuristic poisoning attacks (Carlini and Terzis, 2021; Carlini, 2021; Muñoz-González et al., 2017; Xiao et al., 2012) are specific to threat models related to either conventional regression or classification, and not directly applicable here. The second is that the attack, as poisoning attacks in general, solves a bi-level optimization problem. Principled solutions to this problem, such as Mei and Zhu (2015), make use of the implicit function theorem. However, the approach requires computing an inverse of a Hessian with respect to model parameters, which scales poorly as the model (e.g., neural network) becomes large. Third, prior approaches often assume that learning is convex; in the case of neural network poisoning, however, the problem is non-convex and initialization plays a significant role in attack efficacy.
@@
To address these challenges, we develop a multi-method approach, consisting of two attack methods. The first is a gradient approach in which we relax preference labels to be real-valued, take a series of gradient steps (making use of the implicit function theorem), and project to feasible label space. To deal with unknown initialization seeds, we additionally train attacks to be explicitly robust to a collection of random model initializations, in the spirit of the expectation-over-transformations framework (Athalye et al., 2018) (though the details differ). To handle scalability issues, we couple this approach with either a low-dimensional input embedding learned by training the reward model on clean data, or PCA to reduce input dimensionality. Our second approach relies on a class of rank-by-distance (RBD) greedy heuristics, in which we poison datapoints most similar to the target outcomes the adversary aims to either promote or demote. We instantiate members in this class by defining specific distance functions, such as the Euclidean norm (in input or embedding space) and outcome reward difference.
@@
To understand how preference poisoning manifests in different application domains, we conducted empirical investigation of this vulnerability in four key application domains. The first two domains involve value alignment in control, using MuJoCo simulation environment, which inputs are low-dimensional, and Atari vision-based control setting, in which inputs are images. The third domain uses a recommendation dataset with text as input. Our forth domain involves reward model learning for safety alignment in LLMs, using the LLaMA-7B model (Touvron et al., 2023). We show that the success rate of the best attack varies considerably by domain and even environment within a domain. In many cases, however, our attacks achieve nearly 100% success rate with only 1-10% of data poisoned, and in the case of safety alignment, poisoning only 0.3% of the inputs yields a nearly 100% success rate. A significant observation we make is that which attack is best also varies a great deal across domains: in some cases, the gradient-based approach is best, while in others, RBD methods are better. This demonstrates the importance of our comprehensive multi-method approach to vulnerability analysis. Our final exploration involves several state-of-the-art defense approaches proposed for poisoning attacks. We show that none of these are consistently effective against our attacks. For example, in the case of safety alignment, our attack remains nearly 100% effective even after the application of the defenses.
2. Related Work
There are a wide variation of threat models that fall under the broad category of data poisoning, in which training data is modified to achieve malicious ends (Vorobeychik and Kantarcioglu, 2018). The vast majority are in the supervised learning settings, of which ours can be viewed as a non-conventional variation. We group this body of work into two subcategories: data poisoning and Trojan, or backdoor, attacks.
Data Poisoning Attacks and Defenses For data poisoning attacks and defenses, much work has been focusing on adding new poisoned data, most commonly in the context of classification and regression problems (Biggio et al., 2012, 2011; Jagielski et al., 2018; Jha et al., 2023; Liu et al., 2017a; Malek Esmaeili et al., 2021; Mei and Zhu, 2015; Paudice et al., 2019; Rosenfeld et al., 2020; Steinhardt et al., 2017; Suya et al., 2023b; Wang et al., 2021; Xiao et al., 2012; Yang et al., 2017; Yerlikaya and Bahtiyar, 2022; Zeng et al., 2023). This is quite distinct from our threat model, in which preference responses (labels) are modified, but no new data is added.
The most closely related literature involves label-flipping attacks and defenses in classification (Jha et al., 2023; Malek Esmaeili et al., 2021; Paudice et al., 2019; Rosenfeld et al., 2020; Wang et al., 2021; Xiao et al., 2012; Zeng et al., 2023; Zhang et al., 2020b). Xiao et al. (2012) proposed an optimization framework for identifying the label flips that maximize the classification error, thereby formulating an algorithm for attacking support vector machines. Zhang et al. (2020b) discussed label flipping attacks and defenses for graph neural networks, while Malek Esmaeili et al. (2021) presents an adaptation of differential privacy techniques for defense against label-flipping attacks. Wang et al. (2021) provides a theoretical guarantee of how training (finite-width) networks with stochastic gradient descent can be robust against data poisoning attacks in which labels are flip with a small probability, rather than maliciously selected. Paudice et al. (2019) proposed an algorithm for label-flipping attacks, and a kNN-based algorithm to detect and relabel suspicious data points as a defense. Zeng et al. (2023) proposed a defense approach for several poisoning attacks, including label flipping, that aims to identify a poison-free training data subset. Steinhardt et al. (2017) consider the problem of certified defense against data poisoning, where the defenders first perform outlier removal followed by empirical risk minimization. In most of prior work on label poisoning attacks and defenses, the threat model involves maximizing the prediction error of the model learned on poisoned data. In addition to the distinction in the learning problem structure, this also qualitatively differs from the targeted nature of the promotion and demotion attacks we explore, which is particularly pertinent in the preference learning context. Furthermore, a host of recent approaches for defending against label flipping attack on federated learning have been proposed (Jebreel et al., 2022; Li et al., 2021a; Jebreel et al., 2024; Li et al., 2021b; Awan et al., 2021; Jiang et al., 2023; Manna et al., 2021). Our focus, however, is on the centralized learning paradigm.
Intimately related to our work are poisoning attacks and defenses on recommendation systems (Jia et al., 2023; Fang et al., 2018, 2020; Li et al., 2016; Zhang et al., 2020a). Many target specifically matrix-based or graph-based recommendation systems (i.e., collaborative filtering techniques) (Fang et al., 2018, 2020; Li et al., 2016), which are also distinct from the structure of the MLE reward model learning problem. Most similar to our work is Zhang et al. (2020a), who consider malicious injection of a limited collection of users to poison product ratings so as to promote particular products as a top-K recommendation for the largest number of users. The goal, thus, closely echoes ours, but the learning algorithm being poisoned is quite different.
In poisoning attacks on large language models (LLMs), existing work has focused on text classification tasks (Wallace et al., 2020; Kurita et al., 2020), or achieving specific objectives during instruction tuning (Shu et al., 2023; Wan et al., 2023; Yan et al., 2023). However, these studies have not systematically examined the impact of preference poisoning on reward models, but only on downstream tasks.
Trojan Data Poisoning Attack and Defense Trojan (backdoor) attacks typically entail a combination of data poisoning and implementing a trigger at inference time (Chen et al., 2017; Liu et al., 2020; Boloor et al., 2021; Barni et al., 2019; Li et al., 2020; Baracaldo et al., 2018; Saha et al., 2020; Quiring and Rieck, 2020; Li et al., 2021d; Doan et al., 2021; Nguyen and Tran, 2021; Zhao and Lao, 2022; Salem et al., 2022; Chen et al., 2023). As such, these are inherently distinct as threat models from our setting. An early and prototypical example is due to Gu et al. (2019), who detail how a Trigger can be embedded by adding patterns to inputs and flipping labels to the target. A number of Trojan attacks have also been developed that only modify training labels (in addition to adding a trigger at inference time) (Ge et al., 2021; Chen et al., 2022; Jha et al., 2023). Since the goal in this remains to embed a Trojan which is activated at inference time by adding a Trojan trigger, these remain in nature distinct from our setting. A parallel literature has also developed focusing on defense against Trojan attacks  (Wang et al., 2022a; Cinà et al., 2023). Model Verification is a defense mechanism that checks for anomalies in a model’s functionality to detect potential Trojans (Baluta et al., 2019; He et al., 2019; Erichson et al., 2021; Huster and Ekwedike, 2021). Trojan Trigger Detection aims to detect Trojans by identifying triggers in inputs (Shen et al., 2023; Baracaldo et al., 2018; Liu et al., 2019; Chen et al., 2019; Chakarov et al., 2016; Gao et al., 2019; Xiang et al., 2019; Kolouri et al., 2019; Xu et al., 2019; Huang et al., 2019; Wang et al., 2020; Yoshida and Fujino, 2020). Restoring Compromised Models involves retraining and pruning the neural network or pre-processing to remove triggers (Zhong et al., 2020; Liu et al., 2017b; Zhao et al., 2020; Liu et al., 2018; Wu and Wang, 2021; Zheng et al., 2021; Wang et al., 2019; Guo et al., 2019; Qiao et al., 2019; Chen et al., 2018a; Shen et al., 2021; Veldanda et al., 2020; Udeshi et al., 2019; Villarreal-Vasquez and Bhargava, 2020; Zeng et al., 2020; Du et al., 2019a; Hong et al., 2020). Input filtering aims to cleanse data from malicious inputs, applicable at both training and testing stages (Liu et al., 2022; Wang et al., 2022b; Do et al., 2022; Tran et al., 2018a; Chen et al., 2018b; Tang et al., 2019; Soremekun et al., 2020; Chou et al., 2018; Li et al., 2021c; Fu et al., 2022; Subedar et al., 2019; Du et al., 2019b; Javaheripi et al., 2020; Zeng et al., 2023). None of these, however, are directly pertinent to our problem setting, since we only consider attacks on preference data with no inference-time input modification.
@@ 2. Related Work
3. Learning a Reward Model from Pairwise Comparisons
@@
The problem of learning a reward (utility) model from pairwise comparisons is commonly formalized as follows. Let $D=\{(x_{i},y_{i},o_{i})\}_{i=1}^{n}$ be a dataset of $n$ datapoints with $x_{i},y_{i}\in\mathbb{R}^{m}$ feature vectors representing the $i$th pair of outcomes and $o_{i}\in\{0,0.5,1\}$ a preference between these. Specifically, $o_{i}=0$ if $x_{i}$ is preferred to $y_{i}$ (which we write as $x_{i}\succ y_{i}$), $o_{i}=1$ if $y_{i}\succ x_{i}$, and $o_{i}=0$ if these are preference-equivalent (with $x_{i}\sim y_{i}$ representing such indifference). Typically, data of this kind of obtained by presenting people with pairs of options $(x,y)$ and asking which of these they prefer. For example, this is the process used in improving the helpfulness, and reducing harmfulness, of LLM-based conversational assistants, as part of an RLHF framework (Christiano et al., 2017; Ouyang et al., 2022). Fundamentally, this is a classic problem in utility function learning (Xia, 2019) within the random utility model (RUM) theoretical framework, where one assumed that an agent endowed with a true but unknown utility function $R(x)$ reports preference comparison results corrupted with noise, and one aims to approximately recover the underlying utility function from such data.
@@
(2) $\mathcal{L}(D;\theta)=\sum_{i}-[(1-o_{i})\log\Pr\{x_{i}\succ y_{i}|R_{\theta}\}+o_{i}\log\Pr\{y_{i}\succ x_{i}|R_{\theta}\}].$
4. Threat Model

### v2

@@ Abstract
Learning reward models from pairwise comparisons is a fundamental component in a number of domains, including autonomous control, conversational agents, and recommendation systems, as part of a broad goal of aligning automated decisions with user preferences. These approaches entail collecting preference information from people, with feedback often provided anonymously. Since preferences are subjective, there is no gold standard to compare against; yet, reliance of high-impact systems on preference learning creates a strong motivation for malicious actors to skew data collected in this fashion to their ends. We investigate the nature and extent of this vulnerability by considering an attacker who can flip a small subset of preference comparisons to either promote or demote a target outcome. We propose two classes of algorithmic approaches for these attacks: a gradient-based framework, and several variants of rank-by-distance methods. Next, we evaluate the efficacy of best attacks in both these classes in successfully achieving malicious goals on datasets from three domains: autonomous control, recommendation system, and textual prompt-response preference learning. We find that the best attacks are often highly successful, achieving in the most extreme case 100% success rate with only 0.3% of the data poisoned. However, which attack is best can vary significantly across domains. In addition, we observe that the simpler and more scalable rank-by-distance approaches are often competitive with, and on occasion significantly outperform, gradient-based methods. Finally, we show that state-of-the-art defenses against other classes of poisoning attacks exhibit limited efficacy in our setting.
1 Introduction
@@
Remarkable advances in AI ranging from autonomous systems, such as self-driving cars [3], to conversational agents [97], have led to significant concerns about the possibility that the resulting technologies would come to systematically deviate from conventional norms.
Figure 1: Illustration of Poisoning Attack (Promotion).
@@
Such deviations can range from benign, for example, an occasionally unhelpful response to a prompt, to harmful, such as generating autonomous trajectories that make users (e.g., riders in self-driving cars) fear for their lives or the lives of others, or inciting violence in response to conversational prompts. The fundamental concern is to ensure that the behavior of AI systems aligns with the values of the human communities impacted by these. This value alignment problem has been a subject of foundational research in a broad array of domains, which in addition to autonomous control [16] and conversational agents [62], include recommendation systems [43, 65], and, classically, elections and policymaking [61]. A common foundational element of value alignment, with roots in social choice and utility theory, is to learn a reward model from a dataset comprising human responses about their relative preferences between pairs of options (outcomes, candidates). The learned reward model can then be used in the relevant downstream tasks, such as to make personalized recommendations, or as part of the reinforcement learning from human feedback (RLHF) which has gained popularity in the design of helpful and harmless language models [62, 80, 78].
@@
The nature of preference elicitation necessary in reward model learning, however, opens the door to malicious tampering, as one commonly obtains such information from anonymous subjects. Moreover, the subjective nature of preferences, which often represent a diversity of opinions, makes it very challenging to evaluate the veracity of responses with respect to any gold standard. For example, malicious actors may have an incentive to create fake user accounts, or compromise legitimate users, and thereby populate pairwise comparison datasets collected with preferences responses that achieve their own ends. Consequently, a crucial and under-explored issue is one of vulnerability of reward model learning from pairwise comparison data to malicious preference label poisoning. We aim to systematically investigate this vulnerability within the most common Bradley-Terry (BT) reward model learning framework based on maximum likelihood estimation (MLE) [7, 98], which is a typical approach in RLHF [62, 80, 78].
@@
Our work contributes to the literature on poisoning attacks on machine learning generally [5, 27, 34, 52, 59, 77, 94, 102, 103, 105], and label-flipping attacks in particular [39, 57, 63, 68, 94, 99, 105, 108]. The focus in the existing literature has been predominantly on either classification or regression problems; preference poisoning attacks on reward model learning have received far less attention. Moreover, the problem setting motivates attack goals that focus specifically on relative preferences over feature vectors, which are in themselves distinct from common threat models in prior literature. We show that as a consequence of structural differences in the learning problem and threat model, common defenses against label flipping attacks on conventional supervised learning methods have only limited efficacy in our setting.
@@
There are several technical challenges in devising preference poisoning attacks on reward model learning. First, pairwise preferences are discrete: either one prefers the first option, the second, or is indifferent. Consequently, gradient-based methods proposed for data poisoning, which assume that data is real-valued, cannot be used directly. On the other hand, the typical heuristic poisoning attacks [10, 8, 60, 99] are specific to threat models related to either conventional regression or classification, and not directly applicable here. The second is that the attack, as poisoning attacks in general, solves a bi-level optimization problem. Principled solutions to this problem, such as [59], make use of the implicit function theorem. However, the approach requires computing an inverse of a Hessian with respect to model parameters, which scales poorly as the model (e.g., neural network) becomes large. Third, prior approaches often assume that learning is convex; in the case of neural network poisoning, however, the problem is non-convex and initialization plays a significant role in attack efficacy.
@@
To address these challenges, we propose two classes of attacks. At the core of the first is a projected gradient ascent algorithm that combines iterative relaxation and projection steps with implicit gradient computation. To handle scalability issues, we couple this approach with a low-dimensional input embedding either learned by training the reward model on clean data, or using PCA. The second is a class of rank-by-distance (RBD) greedy heuristics, in which we poison datapoints most similar to the target outcomes the adversary aims to either promote or demote. We instantiate members in this class by defining specific distance functions, such as the Euclidean norm (in input or embedding space) and outcome reward difference.
@@
We conducted empirical investigation of this vulnerability in four key application domains. Our first domain involves reward model learning for safety alignment in LLMs, using the LLaMA-7B model [83]. The next two domains involve value alignment in control, using MuJoCo simulation environment, which inputs are low-dimensional, and Atari vision-based control setting, in which inputs are images. The fourth domain uses a recommendation dataset with text as input. We show that the success rate of the best attack varies considerably by domain and even environment within a domain. In many cases, however, our attacks achieve nearly 100% success rate with only 1-10% of data poisoned, and in the case of safety alignment, poisoning only 0.3% of the inputs yields a nearly 100% success rate. Moreover, in this domain, we show that our poisoning attack has a significant downstream impact on policies learned using RLHF. We also observe that which attack is best varies across domains: in some cases, the gradient-based approach is best, while in others, RBD methods are better. This demonstrates the importance of our comprehensive multi-method approach to vulnerability analysis. Our final exploration involves several state-of-the-art defense approaches proposed for poisoning attacks. We show that none of these are consistently effective against our attacks. For example, in the case of safety alignment, our attack remains nearly 100% effective even after the application of the defenses.
In summary, we make the following contributions:
  1. 1.
@@
    A novel model of poisoning attacks on reward model learning from pairwise comparisons, in which an attacker can flip preference labels.
  2. 2.
    A novel projected gradient ascent approach to solve the combinatorial label poisoning problem.
  3. 3.
    A novel class of rank-by-distance (RBD) poisoning methods that exhibit strong empirical performance, particularly in high-dimensional settings.
  4. 4.
    A theoretical analysis that demonstrates out-of-sample effectiveness of the attacks.
  5. 5.
    Extensive evaluation of the attack efficacy on applications ranging from LLMs to control.
Related Work
There are a wide variation of threat models that fall under the broad category of data poisoning, in which training data is modified to achieve malicious ends [87]. The vast majority are in the supervised learning settings, of which ours can be viewed as a non-conventional variation. We group this body of work into two subcategories: data poisoning and Trojan attacks.
Data Poisoning Attacks and Defenses: Much work on data poisoning attacks and defenses is focused on adding new poisoned data, typically in classification and regression problems [5, 27, 34, 52, 59, 77, 94, 102, 103, 105]. This is distinct from our threat model, in which preference responses are modified, but no new data is added.
The most closely related literature involves label-flipping attacks and defenses in classification [4, 57, 63, 68, 94, 99, 105, 108]. The earlier efforts focused on principled optimization approaches attacking classical machine learning algorithms, such as SVM [4, 99]. More recent efforts, such as [92], which target neural-network-based classifiers, make heavy use of problem-specific heuristics. In most of this work, the threat model involves maximizing the prediction error of the model learned on poisoned data. In addition to the distinction in the learning problem structure, this also qualitatively differs from the targeted nature of the promotion and demotion attacks we explore, which is particularly pertinent in the preference learning context. Moreover, while defense approaches exist, they tend to focus on classification problems that have sufficient structural differences so as to make them insufficient in our setting (see Section 6[57, 63, 77, 105]. Furthermore, while many recent approaches have emerged for defending against label flipping attack on federated learning [37, 48, 38, 49, 2, 42, 58], they are specific to this setting, whereas our focus is on the centralized learning paradigm.
Intimately related to our work are poisoning attacks and defenses on recommendation systems [41, 24, 23, 47, 107]. Many target specifically matrix-based or graph-based recommendation systems (i.e., collaborative filtering techniques) [24, 23, 47], which are also distinct from the structure of the MLE reward model learning problem. Most similar to our work is [107], who consider malicious injection of a limited collection of users to poison product ratings so as to promote particular products as a top-K recommendation for the largest number of users. The goal, thus, closely echoes ours, but the learning algorithm being poisoned is quite different.
In poisoning attacks on large language models (LLMs), existing work has focused on text classification tasks [88, 46], achieving specific objectives during instruction tuning [74, 89, 101], and prompt manipulation in preference data [73, 67]. However, these studies have not systematically examined the impact of preference label poisoning on reward models One recent exception is [92], who study attacks on RLHF by modifying preference labels. Unlike this work, which is restricted to LLMs, our threat model involves general reward model poisoning attacks.
Trojan Attacks and Defenses: Trojan attacks involve a combination of data poisoning and implementing a trigger at inference time [31, 6, 51, 69, 66, 19, 109, 70, 14]. As such, these are inherently distinct as threat models from our setting. For example, [28] detail how a Trigger can be embedded by adding patterns to inputs and flipping labels to the target. A number of Trojan attacks have also been developed that only modify training labels (in addition to adding a trigger at inference time) [26, 13, 39]. A parallel literature has also developed focusing on defense against Trojan attacks  [91, 17]. Prominent examples include model verification, which checks for anomalies in a model’s functionality [29, 22, 33], Trojan trigger detection, which aims to identify triggers [71, 55, 12, 11, 45, 100, 32, 93, 104], restoring compromised models which retrains, prunes, or preprocesses the neural network to remove triggers [113, 56, 110, 54, 96, 112, 90, 72, 85, 86, 106, 21, 30], and input filtering which aims to cleanse data from malicious inputs [53, 95, 18, 82, 76, 15, 50, 25, 20, 36, 105]. None of these, however, are directly pertinent to our problem setting, since we only consider attacks on preference data.
2 Background: Learning a Reward Model from Pairwise Comparisons
@@
The problem of learning a reward (utility) model from pairwise comparisons is commonly formalized as follows. Let $D=\{(x_{i},y_{i},o_{i})\}_{i=1}^{n}$ be a dataset of $n$ datapoints with $x_{i},y_{i}\in\mathbb{R}^{m}$ feature vectors representing the $i$th pair of outcomes and $o_{i}\in\{0,0.5,1\}$ a preference between these. Specifically, $o_{i}=0$ if $x_{i}$ is preferred to $y_{i}$ (which we write as $x_{i}\succ y_{i}$), $o_{i}=1$ if $y_{i}\succ x_{i}$, and $o_{i}=0$ if these are preference-equivalent (with $x_{i}\sim y_{i}$ representing such indifference). Typically, data of this kind of obtained by presenting people with pairs of options $(x,y)$ and asking which of these they prefer. For example, this is the process used in improving the helpfulness, and reducing harmfulness, of LLM-based conversational assistants, as part of an RLHF framework [16, 62]. Fundamentally, this is a classic problem in utility function learning [98] within the random utility model (RUM) theoretical framework, where one assumed that an agent endowed with a true but unknown utility function $R(x)$ reports preference comparison results corrupted with noise, and one aims to approximately recover the underlying utility function from such data.
@@
$\begin{split}\mathcal{L}(D;\theta)=\sum_{i}-[(1&-o_{i})\log\Pr\{x_{i}\succ y_{i}|R_{\theta}\}\\ &+o_{i}\log\Pr\{y_{i}\succ x_{i}|R_{\theta}\}].\end{split}$
Commonly, this loss is minimized using gradient descent with respect to reward model parameters $\theta$.
3 Threat Model